40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 16 02:11:40 2014, Jenne, Update, LSC, CARM and DARM on IR signals, boosts engaged 
    Reply  Wed Apr 16 02:29:30 2014, ericq, Update, LSC, CARM and DARM on IR signals, boosts engaged carmOLTF.pdfboostPlot.png
       Reply  Thu Apr 17 00:49:06 2014, Jenne, Update, LSC, CARM and DARM on IR signals, boosts engaged Lockloss_DARMgainTooHigh_119am.pngLockloss_DARMgainTooHigh_127am.pngSmoothCARMandDARMtransitions_LSCboosts.png
Message ID: 9817     Entry time: Wed Apr 16 02:11:40 2014     Reply to this: 9818
Author: Jenne 
Type: Update 
Category: LSC 
Subject: CARM and DARM on IR signals, boosts engaged 

[Jenne, EricQ]

Tonight, we transitioned CARM and DARM to IR signals, took loop transfer functions, and determined that we could engage the LSC boosts (FM4 in the CARM and DARM servos, which are the same as the XARM and YARM servos). 

Q is preparing spectra to post, and I will dig out time series.  Look for these tomorrow, if they aren't posted tonight.

For the time series data fetching, I have taken notes on what we were doing when, so that I can actually find the data.


11:09pm:  CARM's LSC boost on for the first time

11:14pm:  DARM transferred to AS55Q

11:21pm:  DARM's LSC boost on for the first time

(lockloss)

11:53pm:  CARM transition

12:02am:  DARM transition done, both LSC boosts on

12:04am:  lockloss after reducing CARM digital offset to 0.4

12:45am: PRMI + 2 arms flashing, with no CARM or DARM offsets (arms still on ALS) because we forgot to put in the CARM offset before restoring PRM alignment.  PRMI may have been actually locked, or we may just have been flashing....need to look through the data to see what our recycling looked like.

(lockloss)

1:05am:  pretty smooth transition completed (both CARM and DARM), but we lost lock while reducing the CARM offset.

1:19am: lockloss - why?? We were just sitting at a CARM offset of about 1.3nm (1.3 counts), holding on IR signals.  We were not touching any IFO things while looking at some plots, and just lost lock.  Want to see if we can understand why.

1:27am:  another nice smooth transition for both CARM and DARM to IR signals, but almost immediate lockloss when reducing the CARM offset.


Using the new ALS lock acquisition scripts (elog 9816) and our transition scripts, getting back to PRFPMI lock is pretty smooth and procedural.

* Align arms using ASS (ifo configure screen, restore xarm and yarm, run both arms' ass scripts).

* Align PRMI, no arms (ifo configure screen, restore prmi sideband)

* Find ALS beatnotes, with arm lasers on opposite sides of the PSL.  For both, when increasing the value of the temperature slider, the beatnote should increase in frequency.  (ifo configure screen, restore CARM and DARM als)

* Run ...../scripts/ALS/Lock_ALS_CARM_and_DARM.py

* Run "Find resonance" scripts from ALS screen for each arm.

* Put in a 3 count offset to CARM loop.

* Restore PRM alignment.  (PRMI should acquire lock immediately, although PRM may need some small alignment tweaking).  Enable PRCL and MICH outputs, PRM and BS actuation outputs.

* Reduce CARM offset to 2 counts. 

* Set offsets of 1/sqrt(TRX) and 1/sqrt(TRY) filter banks in the AUXERR section of the LSC screen.  The outputs of both should equal 2 counts (to match the 2 count offset in the CARM loop). 

* Run .../scripts/PRFPMI/Transition_CARM_ALS_to_TransSqrtInv.py , making sure to reduce the CARM digital offset if needed, to keep the arm transmissions at about 0.1 counts.

* Engage FM4 of the CARM filter bank, which is the LSC boost.

* Run .../scripts/PRFPMI/Transition_DARM_ALS_to_AS55.py , making sure to reduce the CARM (or should be DARM?) digital offset if needed, to keep the arm transmissions at about 0.1 counts.

* Engage FM4 of the DARM filter bank, which is the LSC boost.


Notes for going forward:

When we have small-ish digital CARM offsets, such that both of our arm transmitted powers are about 0.1 or higher, we see clear coherence between our CARM_IN1 (the 1/sqrt(trans) signals) and a normalized REFL11_I (again using a spare filter bank like XARM to get REFL11 normalized by (TRX+TRY) ).  We have not yet tried transitioning the CARM digital error signal to this normalized REFL11.

Even though we see that the IFO is much less noisy (as measured by significantly reduced RIN in TRX and TRY as visible by eye on Dataveiwer), we are still losing lock when we reduce the CARM offset.  I have noted above several times, for when we had locklosses, so that I can see if I see anything elucidating in the time series data.

ELOG V3.1.3-